OpenStudy (anonymous):

How many units apart are the points (4, -3) and (-4, 3)?

OpenStudy (anonymous):

Use: \[\large Distance = \sqrt{(x_2 - x_1)^2+ (y_2 - y_1)^2}\] \[\large x_2 = -4, x_1 = 4, y_2 = 3, y_1 = -3\] Plug in the values and find the distance..

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

You are not taking care of negative signs.. \[Distance = \sqrt{(-4-4)^2 + (3 - (-3))^2}\]

OpenStudy (anonymous):

-4 -4 = -8 and: 3 - (-3) = 3 + 3 = 6

OpenStudy (anonymous):

so 2 ?

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

I think I have put square above the brackets in the formula. You have to square them first and then add them..

OpenStudy (anonymous):

i add them together?

OpenStudy (anonymous):

Before adding please square them that is what Formula is saying..

OpenStudy (anonymous):

1 got 100

OpenStudy (anonymous):

Yes right.. Now take the square root of 100 ??

OpenStudy (anonymous):

50

OpenStudy (anonymous):

\(\sqrt{100} = 50 ??\)

OpenStudy (anonymous):

Do you know what is square root ??

OpenStudy (anonymous):

im rite?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Yes now you are right.. Well Done . That is the distance between them..